Any help with any of these would be great thanks

The local department store recently had a large sale event. The store is attempting to figure out the cumulative markdown %. The store took total markdowns of $567,899 and had sales of $3,456,788.

What is the cumulative markdown %?

3.         You are the buyer for a large retail chain. You need to calculate your gross margin. You have the following information available:

Purchases at retail                               $546,843

Purchases at cost                                 324,756

Markdowns 151,324

Employee discounts                                33,456

Freight 4,521

Shortage 1,346

Cash discounts                                          5,678

Sales 359,818

Alterations                                                   342

Senior citizen discount 899

What is the total gross margin?

4.         Repeat exercise three; however, calculate the gross margin using percentages.

5.         A bar is calculating its GM%. You have the following information available, stated in percent to sales:

Markdown % 6.7%

Cash Discounts 1.7%

Shortage 2.6%

Freight 1.4%

Clergy discount                        0.0%

Employee Discount 3.5%

Cumulative Markup % 45.7%

Initial Markup Plan %               46.9%

Advertising Expense %          5.7%

What is the gross margin %?

6.         What is the cumulative markdown % for the bar?

Suppose that after you have finished your calculation the manager says, “Oh, I forgot to give you the percentage for free birthday beers. It was .8% of sales. Would your GM calculation change and, if so, what would the new GM% be?
Answer #1
Markdown percentage = Markdown /Retail Sales
Cumulative Markdown percentage = 567899/(3456788+567899)*100 = 14.11%
3) Gross Margin = Sales - Cost of goods sold
Sales        3,59,818
Beginning stock 0
Purchase at cost 324756
Freight (Direct cost as assumed) 4521        3,29,277
Gross Margin            30,541
4) Gross Margin % = Gross Margin /Sales
Gross Margin % = 30541/359818 = 8.49%